a 1m long resistance wire ,an ammeter and a 2.0v cell are connected in series (the ammeter and the cell have negligible resistance) the ammeter reads 0.50A,calculate,the resistance of the wire?b)the resistance wire is replaced by a 1m long wire of the same material but twice the cross- sectional area ,what is ,i) the resistance of this wire?ii)what is the new ammeters reading?
1
Expert's answer
2018-08-22T11:43:08-0400
R=U/I=2.0/0.50=4 Ω




b) the resistance wire is replaced by a 1m long wire of the same material but twice the cross- sectional area ,what is , i) the resistance of this wire?

Answer:

R_1=R/2=4/2=2 Ω




ii) what is the new ammeters reading?

Answer:

I_1=U/R_1 =2.0/2=1 A